# Machine Learning in Fundamental Physics Natural Image Classification via Quasi-Cyclic Graph Ensembles and Random-Bond Ising Models at the Nishimori Temperature

### v1

@@ Conclusion
By Lemma A.1, Proposition 1, Theorem A.2 and Corollary A.2.1 we obtain the exact correspondence
@@
Let $q$ denote the number of negative eigenvalues of $H_{\beta_{N},J}$. Theorem A.2 shows that each such eigenvalue is in one-to-one correspondence with an independent cycle (i.e. a trapping set) in the graph. Deleting every induced subgraph whose edge set contributes to $q$ forces the Bethe–Hessian to become positive semidefinite, $H_{\beta_{N},J}\;\succeq\;0.$ Because of the Ihara–Bass identity (9), positivity of $H_{\beta_{N},J}$ implies that its determinant never vanishes; consequently all poles of the zeta function disappear. Moreover, the eigenvectors associated with the negative eigenvalues constitute low-energy directions in the spectral embedding; their removal yields a clean, well-conditioned embedding whose geometry is no longer corrupted by spurious modes, thereby improving class separability. In coding-theoretic language these low-energy modes are precisely the topological defects (even cycles with odd degree nodes) that generate pseudocodewords: spin configurations that lie in wide basins of attraction around genuine codewords. When such pseudo-codewords are present, a Bethe–Hessian based training algorithm can become trapped in these basins, leading to sub-optimal solutions and reduced discriminative power for downstream classifiers. By excising the defect-producing subgraphs we simultaneously achieve
@@
From a different perspective, a TS shows up as a negative mode of $H_{\beta,J}$, this mode creates a wide basin of attraction in the underlying energy landscape. Wide basins are precisely what allows an LDPC ensemble to approach Shannon capacity, because they keep the iterative decoder inside a region of locally stable configurations long enough for message passing to converge. The weight-distribution analysis [36] (and its natural extension to replica-symmetry-breaking [37]) shows that, when the degree-distribution pair satisfies the flatness (stability) condition required for capacity on the binary erasure channel – and, by continuity, also for the AWGN case treated with the Bethe–Hessian – the minimum distance can grow only logarithmically with the block length. Hence an ensemble that contains codewords of very large weight (i.e. a trapping set $\mathrm{TS}(a,0)$ with $a\gg\log n$) cannot be capacity-achieving.
Conversely, ensembles enriched with many pseudo-codeword trapping sets, $\mathrm{TS}(a,b\neq 0)$, possess the low-energy negative modes required for those wide basins and can therefore operate arbitrarily close to capacity. The following balance is therefore essential: Modest-weight codewords ($b=0$) guarantee a non-trivial minimum distance, preventing the code from collapsing into a completely dense set of low-weight words; Non-zero-bias trapping sets ($b\neq 0$) generate the wide basins that make iterative decoding robust and enable capacity achievement.
This trade-off creates a controlled imperfection in the graph: it limits the asymptotic growth of $d_{\min}$ while simultaneously improving the discriminative power of the spectral embedding used for classification.
Exactly the same compromise appears in modern learning systems. In paper  [38] introduced the notion of a robust ensemble (RE) – a dense region of weight configurations that can be accessed by interacting replicas (e.g. focusing Belief Propagation, fBP). Their analysis shows that algorithms which bias the dynamics toward such dense regions avoid narrow, poor-performing minima and instead settle in wide, high-entropy basins. These REs are the algorithmic counterpart of the negative Bethe–Hessian modes we identify as trapping sets.

### v2

@@ Conclusion
By Lemma B.1, Proposition 1, Theorem B.2 and Corollary B.2.1 we obtain the exact correspondence
@@
Let $q$ denote the number of negative eigenvalues of $H_{\beta_{N},J}$. Theorem B.2 shows that each such eigenvalue is in one‑to‑one correspondence with an independent cycle (i.e. a trapping set) in the graph. Deleting every induced subgraph whose edge set contributes to $q$ forces the Bethe–Hessian to become positive semidefinite, $H_{\beta_{N},J}\;\succeq\;0.$ Because of the Ihara–Bass identity (9), positivity of $H_{\beta_{N},J}$ implies that its determinant never vanishes; consequently all poles of the zeta function disappear. Moreover, the eigenvectors associated with the negative eigenvalues constitute low‑energy directions in the spectral embedding; their removal yields a clean, well‑conditioned embedding whose geometry is no longer corrupted by spurious modes, thereby improving class separability. In coding‑theoretic language these low‑energy modes are precisely the topological defects (even cycles with odd degree nodes) that generate pseudocodewords: spin configurations that lie in wide basins of attraction around genuine codewords. When such pseudo-codewords are present, a Bethe–Hessian based training algorithm can become trapped in these basins, leading to sub‑optimal solutions and reduced discriminative power for downstream classifiers. By excising the defect‑producing subgraphs we simultaneously achieve
@@
From a different perspective, a TS shows up as a negative mode of $H_{\beta,J}$, this mode creates a wide basin of attraction in the underlying energy landscape. Wide basins are precisely what allows an LDPC ensemble to approach Shannon capacity, because they keep the iterative decoder inside a region of locally stable configurations long enough for message passing to converge. The weight‑distribution analysis [41] (and its natural extension to replica‑symmetry‑breaking [42]) shows that, when the degree‑distribution pair satisfies the flatness (stability) condition required for capacity on the binary erasure channel – and, by continuity, also for the AWGN case treated with the Bethe–Hessian – the minimum distance can grow only logarithmically with the block length. Hence an ensemble that contains codewords of very large weight (i.e. a trapping set $\mathrm{TS}(a,0)$ with $a\gg\log n$) cannot be capacity‑achieving.
Conversely, ensembles enriched with many pseudo‑codeword trapping sets, $\mathrm{TS}(a,b\neq 0)$, possess the low‑energy negative modes required for those wide basins and can therefore operate arbitrarily close to capacity. The following balance is therefore essential: Modest‑weight codewords ($b=0$) guarantee a non‑trivial minimum distance, preventing the code from collapsing into a completely dense set of low‑weight words; Non‑zero‑bias trapping sets ($b\neq 0$) generate the wide basins that make iterative decoding robust and enable capacity achievement.
This trade‑off creates a controlled imperfection in the graph: it limits the asymptotic growth of $d_{\min}$ while simultaneously improving the discriminative power of the spectral embedding used for classification.
Exactly the same compromise appears in modern learning systems. In paper [43] introduced the notion of a robust ensemble (RE) – a dense region of weight configurations that can be accessed by interacting replicas (e.g. focusing Belief Propagation, fBP). Their analysis shows that algorithms which bias the dynamics toward such dense regions avoid narrow, poor‑performing minima and instead settle in wide, high‑entropy basins. These REs are the algorithmic counterpart of the negative Bethe–Hessian modes we identify as trapping sets.
@@
Список литературы
  • [1] M.E. Tipping, et al., "Probabilistic Principal Component Analysis,"J. R. Stat. Soc. Ser. Stat. Methodol. 61, 611–622 (1999). https://doi.org/10.1111/1467-9868.00196.
  • [2] Y. Jia, et al., "Visual concept learning: combining machine vision and Bayesian generalization on concept hierarchies,"NIPS’13, 2, 1842–1850 (2013). https://dl.acm.org/doi/10.5555/2999792.2999818
  • [3] M. Denil, et al., "Predicting parameters in deep learning,"NIPS’13, 2, 2148–2156 (2013). https://dl.acm.org/doi/10.5555/2999792.2999852
  • [4] G. E. Hinton and R. R. Salakhutdinov, "Reducing the Dimensionality of Data with Neural Networks,"Science, 313, 504-507 (2006). https://doi.org/10.1126/science.1127647
  • [5] Wang, C., et al., "LightCSPNet: A Lightweight Network for Image Classification and Objection Detection,"Int J Comput Intell Syst 16, 46 (2023). https://doi.org/10.1007/s44196-023-00226-5
  • [6] L. Dall’Amico et al., "Nishimori meets Bethe: a spectral method for node classification in sparse weighted graphs"J. Stat. Mech. 2021, 093405 (2021). https://doi.org/10.1088/1742-5468/ac21d3
  • [7] V. S. Usatyuk, D. A. Sapozhnikov, S. I. Egorov, "Enhanced Image Clustering with Random-Bond Ising Models Using LDPC Graph Representations and Nishimori Temperature,"Moscow Univ. Phys. Bull. 79, S647–S665 (2025). https://doi.org/10.3103/S0027134924702102
  • [8] H. Nishimori, "Internal Energy, Specific Heat and Correlation Function of the Bond-Random Ising Model,"Prog. Theor. Phys. 66, 1169–1181 (1981). https://doi.org/10.1143/PTP.66.1169
  • [9] M. Hasenbusch et al., "Multicritical Nishimori point in the phase diagram of the ±J Ising model on a square lattice,"Phys. Rev. E 77, 051115 (2008). https://doi.org/10.1103/PhysRevE.77.051115
  • [10] S. Fortunato, D. Hric, Phys. Rep. 659, 1–44 (2016). https://doi.org/10.1016/j.physrep.2016.09.002
  • [11] I. Biazzo, A. Ramezanpour, "Bethe free energy approximations for disordered quantum systems. "Phys. Rev. E 89, 062137 (2014). https://doi.org/10.1103/PhysRevE.89.062137
  • [12] A. Saade et al., "Spectral clustering of graphs with the Bethe Hessian,"NeurIPS 27 (2014). https://dl.acm.org/doi/10.5555/2968826.2968872
  • [13] R. G. Gallager, "Low-Density Parity-Check Codes,"IEEE Trans. Inf. Theory 8, 21–28 (1962). https://doi.org/10.1109/TIT.1962.1057683
  • [14] R. M. Tanner et al., "A class of group-structured LDPC codes,"ISCTA, 365–370 (2001). [Online]. Available: https://www.researchgate.net/publication/2370505_A_class_of_group-structured_LDPC_codes
  • [15] M. P. C. Fossorier, "Quasicyclic low-density parity-check codes from circulant permutation matrices,"IEEE Trans. Inf. Theory 50, 1788–1793 (2004). https://doi.org/10.1109/TIT.2004.831841
  • [16] R. M. Tanner et al., "LDPC block and convolutional codes based on circulant matrices"IEEE Trans. Inf. Theory 50, 2966–2984 (2004). https://doi.org/10.1109/TIT.2004.838370
  • [17] T. J. Richardson and R. L. Urbanke, "Multi-Edge Type LDPC Codes"IEEE ISIT talk (2002). [Online]. Available: http://wiiau4.free.fr/pdf/Multi-Edge%20Type%20LDPC%20Codes.pdf
  • [18] D. Divsalar et al., IEEE ISIT, 1622–1626 (2005). https://doi.org/10.1109/ISIT.2005.1523615
  • [19] I. Stoica et al., "Chord: A scalable peer-to-peer lookup service for internet applications,"ACM SIGCOMM 31, 149–160 (2001). https://doi.org/10.1145/964723.383071
  • [20] R. Khalitov et al., "Sparse factorization of square matrices with application to neural attention modeling,"Neural Netw. 152, 160–168 (2022). https://doi.org/10.1016/j.neunet.2022.04.014
  • [21] M. Sandler et al., "MobileNetV2: Inverted Residuals and Linear Bottlenecks,"IEEE CVPR, 4510–4520 (2018). https://doi.org/10.1109/CVPR.2018.00474
  • [22] G. G. Kasparov, "The Operator K-Functor and Extensions of C*-ALGEBRAS"Math. USSR-Izv. 16, 513–572 (1981). https://doi.org/10.1070/IM1981v016n03ABEH001320
  • [23] B. Vasić et al., "Trapping set ontology,"Allerton, 1–7 (2009). https://doi.org/10.1109/ALLERTON.2009.5394825
  • [24] T. Tian et al., "Construction of irregular LDPC codes with low error floors,"IEEE ICC 5, 3125–3129 (2003). https://doi.org/10.1109/ICC.2003.1203996
  • [25] T. Tian et al., "Selective avoidance of cycles in irregular LDPC code construction,"Trans. Commun. 52, 1242–1247 (2004). https://doi.org/10.1109/TCOMM.2004.833048
  • [26] H. Bass, "The Ihara-Selberg Zeta Function of a Tree Lattice"Int. J. Math. 3, 717–797 (1992). https://doi.org/10.1142/S0129167X92000357
  • [27] M. Atiyah, "Bordism and Cobordism,"Proc. Cambridge Philos. Soc. 57, 200–208 (1961). https://doi.org/10.1017/S0305004100035064
  • [28] P. Bubenik et al., "Multiparameter Persistent Homology via Generalized Morse Theory,"Fields Instit. Comm., 89, 55-79 (2024). https://doi.org/10.1007/978-3-031-57204-3_4
  • [29] Lek-Heng L., "Hodge Laplacians on Graphs,"SIAM Rev. 62, 685–715 (2020). https://doi.org/10.1137/18M1223101
  • [30] S. P. Novikov, "Algebraic construction and properties of Hermitian analogs of K-theory over rings with involution from the viewpoint of Hamiltonian formalism. applications to differential topology and the theory of characteristic classes."Math. USSR-Izv. 4(3), 479–505 (1970).
    https://doi.org/10.1070/IM1970v004n03ABEH000916
  • [31] B. de Tiliere, "The Z-Dirac and massive Laplacian operators in the Z-invariant Ising model,"Electron. J. Probab. 26, 1–86 (2021). https://doi.org/10.1214/21-EJP601
  • [32] J. Bolte and J. Harrison, "Spectral Statistics for the Dirac Operator on Graphs,"J. Phys. A: Math. Gen. 36, 2747–2769 (2003). https://doi.org/10.1088/0305-4470/36/11/307
  • [33] M. F. Atiyah and I. M. Singer, "The index of elliptic operators on compact manifolds"Bull. Am. Math. Soc. 69, 422–433 (1963). https://doi.org/10.1090/S0002-9904-1963-10957-X
  • [34] A. Dua, et al., "Compactifying fracton stabilizer models,"Phys. Review B, 99(24), 245135, (2019). https://doi.org/10.1103/PhysRevB.99.245135
  • [35] J. Dodziuk, V. Mathai, "Approximating $L_{2}$ Invariants of Amenable Covering Spaces: A Combinatorial Approach,"J. Funct. Anal. 154, 359–378 (1998). https://doi.org/10.1006/jfan.1997.3205
  • [36] D. G. Glynn, "The permanent of a square matrix,"Eur. J. Comb. 31, 1887–1891 (2010). https://doi.org/10.1016/j.ejc.2010.01.010
  • [37] R. Smarandache, P. O. Vontobel, "Quasi-Cyclic LDPC Codes: Influence of Proto- and Tanner-Graph Structure on Minimum Hamming Distance Upper Bounds,"IEEE Trans. Inf. Theory 58, 585–607 (2012). https://doi.org/10.1109/TIT.2011.2173244
  • [38] R. Smarandache, "Pseudocodewords from Bethe permanents,"IEEE ISIT, 2059–2063 (2013). https://doi.org/10.1109/ISIT.2013.6620588
  • [39] P. O. Vontobel, "The Bethe Permanent of a Nonnegative Matrix,"IEEE Trans. Inf. Theory 59, 1866–1901 (2013). https://doi.org/10.1109/TIT.2012.2227109
  • [40] V. S. Usatyuk, D. A. Sapozhnikov, "RBIM Natural Image Clustering"[Online]. Available: https://github.com/Lcrypto/Classical-and-Quantum-Topology-ML-toric-spherical/tree/main/Topology_Optim_RBIM_Trapping_Sets
  • [41] C. Di, R. Urbanke and T. Richardson, "Weight distributions: how deviant can you be? IEEE ISIT, 50- (2001), https://10.1109/ISIT.2001.935913
  • [42] M.A. Shokrollahi, "Capacity-Achieving Sequences,"CSG, 123, (2001), https://doi.org/10.1007/978-1-4613-0165-3_9
  • [43] C. Baldassia et al., "Unreasonable effectiveness of learning neural networks: From accessible states and robust ensembles to basic algorithmic schemes,"PNAS, 113(48), E7655-E7662 (2016), https://www.pnas.org/doi/full/10.1073/pnas.1608103113
  • [44] Y. He et al., "Brainstem inhibitory neurons enhance behavioral feature selectivity by sharpening the tuning of excitatory neurons,"Curr Biol. , 34(20):4623-4638 (2024). https://doi.org/10.1016/j.cub.2024.08.037
  • [45] E. I. Knudsen, "Neural Circuits That Mediate Selective Attention: A Comparative Perspective,"Trends Neurosci., 41(11), 789–805 (2018), https://doi.org/10.1016/j.tins.2018.06.006